A number minus 11 equals 12 minus the number.

Answer by nerdybill(7384)

A number minus 11 equals 12 minus the number.
Let x = the number
then
n - 11 = 12 - n
2n - 11 = 12
2n = 23
n = 23/2
n = 11.5